# DEGAP: Dual Event-Guided Adaptive Prefixes for Templated-Based Event Argument Extraction Model with Slot Querying

### v1

@@ B.2 Baselines
  • TSAR Xu et al. (2022): A classification-based model that first identifies candidate spans and then predicts their roles within events. It utilizes an abstract meaning representation (AMR) graph to help learn the better representations of candidate spans. We report the results from He et al. (2023).
  • SCPRG Liu et al. (2023c): A classification-based model that enhances the final candidate span representations using context information filtered by event trigger and role interaction features. We use their code [^7] https://github.com/LWL-cpu/SCPRG-master to test its performance on all datasets.
@@
  • TabEAE He et al. (2023): Building upon PAIE, it simultaneously extracts multiple events and replaces the base model BART used by PAIE with RoBERTa. The original paper introduces three training-inference paradigms: Multi-Multi, Multi-Single, and Single-Single (consistent with our training-inference paradigm). To ensure a fair comparison, we report the performance of its Single-Single version.
B.3 The implementation details of DRAAE
@@
Appendix C Experimental Analysis
C.1 Performance Comparison with LLMs
@@ C.2 Efficiency Analysis
C.3 Event Template Variant

### v2

@@ C.3 Baselines
  • TabEAE He et al. (2023): Building upon PAIE, it simultaneously extracts multiple events and replaces the base model BART used by PAIE with RoBERTa. The original paper introduces three training-inference paradigms: Multi-Multi, Multi-Single, and Single-Single (consistent with our training-inference paradigm). To ensure a fair comparison and visually demonstrate the effectiveness of our proposed prefix, we report the performance of its Single-Single version.
C.4 The implementation details of DRAAE
@@
Appendix D Experimental Analysis
D.1 Performance Comparison with LLMs
@@ D.2 Efficiency Analysis
D.3 Prefix length
We further analyze the impact of prefix length on model performance. Intuitively, a longer prefix implies more trainable parameters, thus enabling it to provide more expressive guidance for the model regarding relevant semantic information. As shown in Fig. LABEL:prefix_len, as the prefix length increases, the model performance tends to improve, followed by a slight decline. We speculate that this is because too long prefixes provide redundant semantic information, making it difficult for the model to focus on extracting the target event.
D.4 Event Template Variant